# Anytime-valid t-tests and confidence sequences for Gaussian means with unknown variance

### v3

@@ A.2 A Sequential Test for Median by Howard and Ramdas [2022]
It is straightforward to see that the growth of $\{B_{n}^{\lambda}\}$ under some $N_{\mu,\sigma^{2}}$ has the following growth index:
@@
Here, $\Phi$ is the cumulative distribution function of $N_{0,1}$. As is shown in Figure 4, this is a very small growth index compared to the optimal $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$, and is powerful only against the alternative $\mathcal{N}_{\mu>0}$ when $\lambda>0$. Therefore, Howard and Ramdas [2022] consider different choices of $\lambda$ for different epochs of time (“stitching”) to obtain Proposition A.4, which is still much looser than many of the other CSs we derive. Another way to gain universal power against all $\mathcal{N}_{\mu\neq 0}$ is, again, via the method of mixture, and Howard and Ramdas [2022] consider a beta-binomial mixture distribution on $\lambda$. Below, we let $\operatorname{B}(a,b)$ denote the beta function.
@@
is a supermartingale on $\{\mathcal{F}_{n}\}_{n\geqslant 0}$ for the set of all continuous distributions with median 0; in particular, for $\mathcal{N}_{\mu=0}$. Under $N_{\mu,\sigma^{2}}$, it has the following growth index:
@@
Figure 4: The growth indices of the one-sided median test supermartingale (78) and mixture test supermartingale (80) under different choices of $\lambda$ and alternative effect sizes $\mu/\sigma$. The “upper bound” refers to the optimal growth index $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$.
This statement, proved in Section C.5, shows that the mixture does gain power over all $\mu\neq 0$ as is depicted in Figure 4. However, compared to the optimal growth index $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$, it falls short in power especially when the effect size $\mu/\sigma$ is large.
@@
The convergence (80) also holds in expectation (because the sequence is uniformly integrable). Following Vovk and Wang [2022, Section 5], we may define a concept of asymptotic relative efficiency (ARE) of the test process $\{B_{n}^{(a,b)}\}$ here, as the ratio between the second derivatives at 0 of the growth index curves of the upper bound and $\{B_{n}^{(a,b)}\}$ in Figure 4:
@@ C.4 Proofs for Optimality
Proof of Proposition 6.4.
@@
This concludes that the likelihood ratio is a supermartingale for $\mathcal{N}_{\mu=0}$. The rest of the statement follows from the basic facts we mentioned about numeraires and the reference therein. ∎
@@ C.5 Proofs for Additional t-Tests
we can see that the test supermartingale $\{B_{n}^{(a,b)}\}$ under some $N_{\mu,\sigma^{2}}$ has the following growth index:

### v4

@@ A.2 A Sequential Test for Median by Howard and Ramdas [2022]
It is straightforward to see that the growth of $\{B_{n}^{\lambda}\}$ under some $N_{\mu,\sigma^{2}}$ has the following e-power:
@@
Here, $\Phi$ is the cumulative distribution function of $N_{0,1}$. As is shown in Figure 4, this is a very small e-power compared to the optimal $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$, and is powerful only against the alternative $\mathcal{N}_{\mu>0}$ when $\lambda>0$. Therefore, Howard and Ramdas [2022] consider different choices of $\lambda$ for different epochs of time (“stitching”) to obtain Proposition A.4, which is still much looser than many of the other CSs we derive. Another way to gain universal power against all $\mathcal{N}_{\mu\neq 0}$ is, again, via the method of mixture, and Howard and Ramdas [2022] consider a beta-binomial mixture distribution on $\lambda$. Below, we let $\operatorname{B}(a,b)$ denote the beta function.
@@
is a supermartingale on $\{\mathcal{F}_{n}\}_{n\geqslant 0}$ for the set of all continuous distributions with median 0; in particular, for $\mathcal{N}_{\mu=0}$. Under $N_{\mu,\sigma^{2}}$, it has the following e-power:
@@
Figure 4: The e-powers of the one-sided median test supermartingale (78) and mixture test supermartingale (80) under different choices of $\lambda$ and alternative effect sizes $\mu/\sigma$. The “upper bound” refers to the optimal e-power $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$.
This statement, proved in Section C.5, shows that the mixture does gain power over all $\mu\neq 0$ as is depicted in Figure 4. However, compared to the optimal e-power $\frac{1}{2}\log(1+\mu^{2}/\sigma^{2})$, it falls short in power especially when the effect size $\mu/\sigma$ is large.
@@
The convergence (80) also holds in expectation (because the sequence is uniformly integrable). Following Vovk and Wang [2024, Section 5], we may define a concept of asymptotic relative efficiency (ARE) of the test process $\{B_{n}^{(a,b)}\}$ here, as the ratio between the second derivatives at 0 of the e-power curves of the upper bound and $\{B_{n}^{(a,b)}\}$ in Figure 4:
@@ C.4 Proofs for Optimality
Discussion on Proposition 6.5.
Before we present and prove an extended version of Proposition 6.5, let us briefly review some key definitions and results of Larsson et al. [2024]. Let $\mathcal{P}$ and $Q$ be mutually equivalent probability measures. Let $\mathcal{P}^{\circ\circ}$ be the “bipolar” of the measure set $\mathcal{P}$, which contains all sub-probability measures for which any e-value for $\mathcal{P}$ is an e-value as well. $\mathcal{P}^{\circ\circ}$ is thus a superset of $\mathcal{P}$, understood as the “effective null”. If the infimum $\inf_{P\in\mathcal{P}^{\circ\circ}}D_{\operatorname{KL}}(Q\|P)$ is attained as minimum by some $P^{*}\in\mathcal{P}^{\circ\circ}$, we say that $P^{*}$ is the reverse information projection of $Q$ onto the set $\mathcal{P}$. Proposition 2.4, Theorem 3.4, and Theorem 4.1 of Larsson et al. [2024] jointly state the following.
Lemma C.2.
Let $P^{*}$ be a sub-probability measure such that $M^{*}=\frac{\mathrm{d}Q}{\mathrm{d}P^{*}}(X_{1})$ is an e-value for $\mathcal{P}$. Then, the following are equivalent:
  1. 1.
    $P^{*}\in\mathcal{P}^{\circ\circ}$;
  2. 2.
    $P^{*}$ is the reverse information projection of $Q$ onto $\mathcal{P}$;
  3. 3.
    $M^{*}$ is the optimal e-value for $\mathcal{P}$: any e-value $M$ for $\mathcal{P}$ satisfies $\mathrm{E}_{Q}(\log M)\leqslant\mathrm{E}_{Q}(\log M^{*})$.
When this happens, the optimal e-value $M^{*}$ is called the numeraire e-value.
Proposition C.3 (Reverse information projection and numeraire in the t-test).
The probability measure $N_{0,\sigma^{2}+\mu^{2}}$ attains the following minimum
$D_{\operatorname{KL}}(N_{\mu,\sigma^{2}}\|N_{0,\sigma^{2}+\mu^{2}})=\min_{P\in\mathcal{N}_{\mu=0}}D_{\operatorname{KL}}(N_{\mu,\sigma^{2}}\|P)=\frac{1}{2}\log(1+\mu^{2}/\sigma^{2}).$ (207)
Further, $M^{*}=\frac{\mathrm{d}N_{\mu,\sigma^{2}}}{\mathrm{d}N_{0,\sigma^{2}+\mu^{2}}}(X_{1})$ is an e-value for $\mathcal{N}_{\mu=0}$, thus it is the numeraire e-value and $N_{0,\sigma^{2}+\mu^{2}}$ the reverse information projection of $N_{\mu,\sigma^{2}}$ onto $\mathcal{N}_{\mu=0}$. Consequently, any e-value $M$ for $\mathcal{N}_{\mu=0}$ has e-power at most
$\mathrm{E}_{N_{\mu,\sigma^{2}}}(\log M_{n})\leqslant\frac{1}{2}\log(1+\mu^{2}/\sigma^{2}),$ (208)
equality if and only if $M$ is the numeraire $M^{*}$.
Proof of Proposition C.3.
@@
This concludes that the likelihood ratio is an e-value for $\mathcal{N}_{\mu=0}$. The rest of the statement follows from Lemma C.2. ∎
@@ C.5 Proofs for Additional t-Tests
we can see that the test supermartingale $\{B_{n}^{(a,b)}\}$ under some $N_{\mu,\sigma^{2}}$ has the following e-power: